Determine the required value of the missing probability to make the
distribution a discrete probability distribution.
x	P(x)
3	0.33
4	?
5	0.23
6	0.17

Found 2 solutions by ikleyn, Edwin McCravy:
Answer by ikleyn(52814)

.

The sum of values in the column called P(x) must be equal to 1.



Answer by Edwin McCravy(20060)

Determine the required value of the missing probability to make the
distribution a discrete probability distribution.
x	P(x)
3	0.33
4	?
5	0.23
6	0.17

--------------------------------
The probabilities have to add up to 1. Put an N for the question mark:

x	P(x)
3	0.33
4	N
5	0.23
6	0.17

Add 'em up, and set it equal to 1:

0.33 + N + 0.23 + 0.17 = 1

              N + 0.73 = 1

                     N = 1 - 0.73

                     N = 0.27

x	P(x)
3	0.33
4	0.27
5	0.23
6	0.17

Now it's a discrete probability distribution.
That's all there is to it!

Edwin